Reports of 19802
VIS (PR3)
takafumi.ushiba - 23:41 Wednesday 10 August 2022 (21754) Print this report
Comment to PR3 OpLev rearangement (21753)

[Yamamoto, Ushiba]

After this work, we found one of the IM OSEM signal cable was disconnected and cannot reach LOCK_ACQUISITION state.
This situation was already recovered but please do not leave the screws of cables loose, or this kind of happening won't go away from KAGRA.

At least on PR3 satellite box, we found three unlocked cables and five loose cables out of 15; it's terrible.

VIS (PR3)
takafumi.ushiba - 23:34 Wednesday 10 August 2022 (21753) Print this report
PR3 OpLev rearangement

[Yokozawa, Yasui, Ushiba]

Abstract:
We rearanged PR3 OpLevs to avoid sticking.out QPDs form Optical tables on the pylon.
Rearangement was succeeded and new wind shield made by aluminum foils were constracted.

Detail:
<Motivation>
Since PR3 OpLev output beam was passing through the very edge of the viewport and QPDs for the tilt OpLev was installed sticking out from the optical table.
So, we moved injection beam position and rearaged optical layout to avoid protrusion.

<What we did>
We moved injection beam position and adjusted the injection angle in order not to clip the beam.
After that, we rearange the reciever side optical components.
To avoid changing calibration factors, we set a lens, mirror, and QPD for length sensing OpLev as follows:
*Distance between BS and lens: 25 mm
*Distance between lens and mirror: 215 mm
*Distance between mirror and QPD: 215 mm
According to the klog19613, total diatance from BS to QPD is 455 mm, which is the same as the current setup.

After that, we constracted wind shield for OpLevs.
In this time, we didn't make windshield between viewport flanges and pylons because we will soon have leak test of central area after vacuum evacuation.
So, we need to make it after cnfirming vacuum leak was not there.
In addition, we made aluminum foil wind shield so as to be easily open by only detaching white tapes.
So, please do not touch pink tapes on the windshield, otherwise the windshield will be broken and need to be repaired from the begining. 

After the work, I recovered the alignment and current setpoint of PR3 is a good enough to lock the X arm.

Comments to this report:
takafumi.ushiba - 23:41 Wednesday 10 August 2022 (21754) Print this report

[Yamamoto, Ushiba]

After this work, we found one of the IM OSEM signal cable was disconnected and cannot reach LOCK_ACQUISITION state.
This situation was already recovered but please do not leave the screws of cables loose, or this kind of happening won't go away from KAGRA.

At least on PR3 satellite box, we found three unlocked cables and five loose cables out of 15; it's terrible.

MIF (ASC)
takafumi.ushiba - 21:52 Wednesday 10 August 2022 (21752) Print this report
Comment to First development of identifying IR center position by TCam photo (21742)

Thank you for your great efforts on the beam position confirmation.
I would like to ask you to apply this analysis to X arm Green beam and Y arm IR beam because current alignment is performed by checking GRX beam position on ETMX and IRY beam position on ETMY.

VIS (General)
ryutaro.takahashi - 15:29 Wednesday 10 August 2022 (21751) Print this report
F1 damper

[Takahashi, Fabian, Hirata]

We took pictures of the gap between the damper ring and the Cu plate in SRM and SR3. The gap in SRM (~3mm) looks smaller than the gap in SR3 (~4mm). The measured transfer functions from the IP yaw actuator to the TM yaw Oplev showed almost the same Qs for the lowest mode. The height of the F0 keystone in the pictures deviated from the height at the TF measurement a little (+40um for SRM and +270um for SR3).

Images attached to this report
AOS (Baffles & Beam dumps)
koji.nakagaki - 15:10 Wednesday 10 August 2022 (21750) Print this report
Comment to Invac work finalization in IFI and IMM chambers: Day 3 (21638)
[Yokozawa,Nakagaki]

The IP address and other settings were configured on the IFI CRYOCON and connected to the network.

IPaddress: 172.16.34.149

MIF (General)
takahiro.yamamoto - 14:04 Wednesday 10 August 2022 (21749) Print this report
Comment to Turn on PZT drivers and add PZT offsets for POP STM1 and STM2 (21730)

It comes from the digital AI process between user models and IOP models.

On real-time system, default implementation of the digital AI process is 0-padding way.
FIg.1 shows a simulated signal on the 65kHz IOP model when a 2kHz user model outputs 10000ct offset.
We can see a periodical (2kHz) glitch like signals.
By decreasing offset value on the user model, an amplitude of the glitch like signal also decrease (See also Fig.2 with 1000ct offset).
Peak amplitude of these glitches seems to become ~27% larger than the given DC offset value.

We have two choices to avoid this noise.

One is to change the model rate from 2kHz to 16kHz.
A simulated signal on the IOP model when 16kHz user model outputs 10000ct offset is shown in Fig.3.
In this case, 16kHz noise is induced but peak amplitude can be suppressed as ~4% of the given DC offset value.
Thanks to the response of cavity, PZT, etc. noise around 16kHz may be practically negligible.

Another choice is to use "no_zero_pad" option on the user model.
In this case, we can completely remove this kind of noise. But CPU load increase.
Fortunately, CPU load is not so serious because k1ial is now very small model.
(It can be used on very large model e.g k1lsc, k1asc, and so on.)

I haven't used this option and seen it also on LIGO's models yet.
So we should test it if we choose the latter choice.

Images attached to this comment
VIS (IX)
takaaki.yokozawa - 11:54 Wednesday 10 August 2022 (21748) Print this report
Comment to Investigation of the glitch in the MN oplev light source : Effect of the fiber collimator (21648)
[Chiba, Yokozawa]

I checked the long term stability of the IX MN YPL oplev SUM.
It seems quite stable compared from previous one.
There are some glitch like signal (like around 2d), there would be actual motion (Some temperature drift or slow SUM change)
Images attached to this comment
MIF (General)
takafumi.ushiba - 11:07 Wednesday 10 August 2022 (21747) Print this report
Comment to Turn on PZT drivers and add PZT offsets for POP STM1 and STM2 (21730)

[Yasui, Ushiba]

We checked DAC outputs from corresponding channels and found that there are strange 2048-Hz signals in the DAC output itself (fig1).
During ths measurement, we disconnected the BNC cables between DGS rack and PZT drivers, so it is not likely to be related GND of the drivers.

Images attached to this comment
CRY (Cryostat IY)
shinji.miyoki - 9:18 Wednesday 10 August 2022 (21745) Print this report
Comment to Turned (21710)

The IYV temp reached 22.6 and fluctuated.

Because I doubted some troubles of the air conditioner, I asked Hayakawa-kun to check it. However, he could not find any unusual situation in it.

Images attached to this comment
VAC (EY)
nobuhiro.kimura - 8:47 Wednesday 10 August 2022 (21746) Print this report
Equipments in Y-end were moved to the center room
When: 9:30~14:30 09/Aug./2022
Who: Kimura, Yasui, M. Takahashi and Uchiyama
Task:
1. Equipments for leak testing and repair metal gaskets in Y-end were moved to the center room.
The Mozumi Tunnel was used as a route for transporting the equipment out.
2. Residual gas measurements were performed on IYC, EYC, and IXC.
We attached results of residual gas measurement for IXC, EXC as a reference.
Images attached to this report
CAL (YPcal)
dan.chen - 6:15 Wednesday 10 August 2022 (21744) Print this report
Integration sphere position marking and optical efficiency measurement

Date: 2022/8/9

With Takashi Takase

We checked beam alignment, and adjusted WSK (old name: GSK) position against the Pcal beams, then marked the positions for higher reproducibility during the integration sphere calibration.

Details: link

MIF (General)
takahiro.yamamoto - 20:57 Tuesday 09 August 2022 (21743) Print this report
Comment to Turn on PZT drivers and add PZT offsets for POP STM1 and STM2 (21730)

[Ushiba, YamaT]

After this work, 2048Hz and its harmonics (4096Hz and 6144Hz) appears on the control signals of Gr PDH-X and ALS DARM as shown in Fig.1.
Blue and red curves represent before (around 9:00 on 8/8 JST) and after (around 15:00 on 8/8) spectra, respectively.
There is no line on 1024Hz and 3072Hz. So it seems to be a series of 2048Hz.

When we took the red curve, PZT output is 75V (~16384ct) as the offset form DAC output.
And these lines can be removed by changing the DAC output as 0ct (0V).
So the output to POP PZT induces these lines.

Because lines lie just 2048, 4096, and 6144Hz with the frequency resolution as 0.125Hz.
So it seems to be caused from digital issue not from PZT itself, circuits etc.

Tomorrow, Ushiba-kun will check the output signals from Dsub-BNC converter to check that
a cause lie around digital side or PZT side just in case.

We set PZT offset as 0 for today.

Images attached to this comment
MIF (ASC)
ayumu.toriyama - 19:09 Tuesday 09 August 2022 (21742) Print this report
First development of identifying IR center position by TCam photo

[Kenta Tanaka, Yuzurihara, Shiota (Aogaku), S. J. Tanaka (Aogaku), Toriyama (Aogaku)]

Because ITMX and ITMY has birefringence, in order to apply ASC, the relative position between mirror center snd IR position shold be known by TCam.

Aogaku member stayed at KAGRA site from Aug. 3 to Aug. 10.

We developed 1st version code to estimate the IR center position by TCam photo.

I attached the example of IX and EX. 

After further code test the code will be included in TCam system.

 

Images attached to this report
Comments to this report:
takafumi.ushiba - 21:52 Wednesday 10 August 2022 (21752) Print this report

Thank you for your great efforts on the beam position confirmation.
I would like to ask you to apply this analysis to X arm Green beam and Y arm IR beam because current alignment is performed by checking GRX beam position on ETMX and IRY beam position on ETMY.

VIS (BS)
ryutaro.takahashi - 18:25 Tuesday 09 August 2022 (21741) Print this report
F1 damper

[Takahashi, Fabian, Hirata]

  • I measured the transfer function from the IP yaw actuator to the TM yaw Oplev in the safe mode. Though it was difficult to measure the TF of the quadruple pendulum, the lowest mode at 43mHz had a Q of about 5 (coherence was about 0.7 at the frequency).
  • We doubt if the F1 damper is working well. We went to take pictures of the gap between the damper ring and the Cu plate with the 360 camera. The gap was about 5mm. We also check the gap in SR2.
  • We found the fishing rod for the F0 was very close to the upper limit. It should be centered by removing the ballast weight on the F0 keystone.
Images attached to this report
VIS (General)
fabian.arellano - 17:19 Tuesday 09 August 2022 (21740) Print this report
Type-B suspension magnet polarities and lower breadboard damper ring position.

With Hirata-san and Takahashi-san.

We measured the magnet polarities in the IPs and the IMs. In all of them:

  • Intermediate Mass:
    • H1, H2, H3, V2, V3: North out
    • V1: South out
  • Inverted pendulum yokes (I'll post a diagram later):
    • Top: North
    • Middle: South
    • Bottom:  North

We also measured the gap between the damper ring of the lower breadboard and the corresponding copper rings (I'll post a diagram later):

  • BS: we forgot
  • SR2: 22 mm
  • SR3: 20 mm
  • SRM: 20 mm
DetChar (General)
shoichi.oshino - 9:14 Tuesday 09 August 2022 (21739) Print this report
SummaryPage trouble: process was stopped from 8/8 7:00 UTC
SummaryPage GIF process was stopped from 8/8 7:00 UTC.
I deleted this process.
After this work, SummaryPage starts to create plots.
DetChar (General)
shoichi.oshino - 9:11 Tuesday 09 August 2022 (21738) Print this report
SummaryPage trouble: process was stopped from 8/7 7:00 UTC
SummaryPage VIS process was stopped from 8/7 7:00 UTC.
I deleted this process.
After this work, SummaryPage starts to create plots.
CRY (Cryostat IY)
shinji.miyoki - 7:54 Tuesday 09 August 2022 (21736) Print this report
Comment to Turned (21710)

The IYV temp keeps gradually increasing.

Images attached to this comment
CAL (XPcal)
dan.chen - 6:16 Tuesday 09 August 2022 (21735) Print this report
Integration sphere position marking and optical efficiency measurement

Date: 2022/8/8

With Takashi Takase

We checked beam alignment (link), and adjusted WSK (old name: GSK) position against the Pcal beams, then marked the positions for higher reproducibility during the integration sphere calibration.

Details: link

CAL (XPcal)
dan.chen - 6:15 Tuesday 09 August 2022 (21734) Print this report
XPcal beam position check

Date: 2022/8/8

I checked XPcal beam positions on ETM, and the positions were in acceptable regions.

Details: link

MIF (General)
kenta.tanaka - 20:28 Monday 08 August 2022 (21732) Print this report
Comment to Turn on PZT drivers and add PZT offsets for POP STM1 and STM2 (21730)

## Abstract

I calculated the beam position displacement on ITM with ABCDEF matrix. According to PZT specsheet, PZT applied at 75V pushes the mirror forward 1.8 um (= 3.6um/150V*75V). When STM1 and STM2 pushed the mirror forward by 1.8um each, the beam spot position on the ITM shifted ~0.1 mm from the position before pushing.

### How to calculate it

  • The propagation of rays in a system slightly misaligned from the ABCD system is calculated with the matrix called "ABCDEF matrix", which is an extension of the original 2x2 ABCD matrix to a 3x3 matrix by adding a shift of the beam position due to misalignment to the elements (1,3), a tilt of the optical axis due to misalignment to the elements (2,3), and (0,0,1) in the third row.
  • The ABCDEF matricies of free space propagation, reflection from a displaced and misaligned mirror and thin lens is represented as below.
    • free propagation L^free=[1l0010001]\hat{L}_{\mathrm{free}} = \left[ \begin{matrix} 1 & l & 0 \\ 0 & 1 & 0 \\ 0 & 0 & 1 \end{matrix} \right]
    • reflection from a mirror R^mirror=[102Δxsinθ-2Re12Δα001]\hat{R}_{\mathrm{mirror}} = \left[ \begin{matrix} 1 & 0 & 2\Delta x \sin \theta \\ -\frac{2}{R_e} & 1 & 2\Delta \alpha \\ 0 & 0 & 1 \end{matrix} \right]
    • thin lens F^lens=[100-1f10001]\hat{F}_{\mathrm{lens}} = \left[ \begin{matrix} 1 & 0 & 0 \\ -\frac{1}{f} & 1 & 0 \\ 0 & 0 & 1 \end{matrix} \right]
      • where l is distance in the free space, Δx\Delta x is lateral shift of a mirror, θ\theta is an incident angle to a mirror, ReR_e  is an effective radious of curvature of a mirror, Δα\Delta \alpha is a misaligned angle of a mirror and ff is a focus length of a lens (c.f. Siegman "Lasers" chap. 15)
  •  The ABCDEF matrix from STM1 to ITM is as below
    •  M^=L^pr3itmR^pr3L^pr2pr3L^pr2GrL3F^GrL3L^GrL3GrM3R^GrM3L^GrM3GrM2R^GrM2\hat{M} = \hat{L}_{\mathrm{pr3itm}} \hat{R}_{\mathrm{pr3}} \hat{L}_{\mathrm{pr2pr3}} \hat{L}_{\mathrm{pr2GrL3}} \hat{F}_{\mathrm{GrL3}} \hat{L}_{\mathrm{GrL3GrM3}} \hat{R}_{\mathrm{GrM3}} \hat{L}_{\mathrm{GrM3GrM2}} \hat{R}_{\mathrm{GrM2}}
  • Assuming that the ray vector before shifting is r0=[00]r_0 = \begin{bmatrix} 0 \\ 0 \end{bmatrix}, and that STM1 and STM2 are shifted forward by 1.8um (= 3.6 um /150 V * 75 V) each and no misalignment has occurred, the beam shift on the ITM is equal to M(1,3).
    • M(1,3) ~ 0.1 mm
    • R_e
MIF (General)
takafumi.ushiba - 19:23 Monday 08 August 2022 (21730) Print this report
Turn on PZT drivers and add PZT offsets for POP STM1 and STM2

Related work with klog21163.

[Toriyama, Ushiba]

Abstract:
We turned on PZT drivers and added 75 V offsets for each PZTs from DGS.
Alignment seems to come back after adding all PZT offset, so we didn't change anything else.

Detail:
I implemented EUL2PZT matrices for STM1 and STM2 as shown in fig1.
Then, I implemented calibration factor from PZT driver output voltage to corresponding DGS cnts in K1:ALS-X_IAL_STM{1,2}_PZT filter banks (fig2).
Since 0V (0 cnts) - 10V (2^15 cnts) are corresponding to 0V - 150V, a calibration factor is (2^15)/150 = 218.45 cnts/V.
Thanks to this calibration factors, we can set the offset of the filter banks as the voltage that we would like to apply to the PZTs.


We connected PZT drivers to AC power supply, which is used for PDA1 of fiber output monitors and prepared inside the POP table.
After that, we turned on PZT drivers and set analog outputs from the driver to zeros.


First, we added offsets for PZTs of STM2 one by one.
Figure 3 shows the time series data of transmitted GRX signals during the work.
As you can see, transmitted light power once go down but come back after all offsets were added.

In the same way, we aded offsets for PZTs of STM1 one by one.
Figure 4 shows the time trend of  transmitted GRX signals during the work.
In this time, we cannot see any significant drop of the signals.

After that, I requested the XARM guardian to ALS_LOCKED state and then we can lock XARM with TEM00 (fig5).
Since the transmitted light power didn't change so much (~1.5: fig6), we didn't perform any additional alignment workafter adding PZT offsets.

Note:
As I wrote, PZT drivers are connected to the same power supply for PDA1 of fiber output monitor now.
I'm not sure it shold be avoided but even if so, we don't have to hesitate changing PZT driver power supply because we confirmed PZTs don't affect the alignment so much if we move all of them at the same time.

Tanaka-kun kindly calculated how much beam spot change is expected due to this work and it seems negligible.
Detail will be reported by Tanaka-kun.

Images attached to this report
Comments to this report:
kenta.tanaka - 20:28 Monday 08 August 2022 (21732) Print this report

## Abstract

I calculated the beam position displacement on ITM with ABCDEF matrix. According to PZT specsheet, PZT applied at 75V pushes the mirror forward 1.8 um (= 3.6um/150V*75V). When STM1 and STM2 pushed the mirror forward by 1.8um each, the beam spot position on the ITM shifted ~0.1 mm from the position before pushing.

### How to calculate it

  • The propagation of rays in a system slightly misaligned from the ABCD system is calculated with the matrix called "ABCDEF matrix", which is an extension of the original 2x2 ABCD matrix to a 3x3 matrix by adding a shift of the beam position due to misalignment to the elements (1,3), a tilt of the optical axis due to misalignment to the elements (2,3), and (0,0,1) in the third row.
  • The ABCDEF matricies of free space propagation, reflection from a displaced and misaligned mirror and thin lens is represented as below.
    • free propagation L^free=[1l0010001]\hat{L}_{\mathrm{free}} = \left[ \begin{matrix} 1 & l & 0 \\ 0 & 1 & 0 \\ 0 & 0 & 1 \end{matrix} \right]
    • reflection from a mirror R^mirror=[102Δxsinθ-2Re12Δα001]\hat{R}_{\mathrm{mirror}} = \left[ \begin{matrix} 1 & 0 & 2\Delta x \sin \theta \\ -\frac{2}{R_e} & 1 & 2\Delta \alpha \\ 0 & 0 & 1 \end{matrix} \right]
    • thin lens F^lens=[100-1f10001]\hat{F}_{\mathrm{lens}} = \left[ \begin{matrix} 1 & 0 & 0 \\ -\frac{1}{f} & 1 & 0 \\ 0 & 0 & 1 \end{matrix} \right]
      • where l is distance in the free space, Δx\Delta x is lateral shift of a mirror, θ\theta is an incident angle to a mirror, ReR_e  is an effective radious of curvature of a mirror, Δα\Delta \alpha is a misaligned angle of a mirror and ff is a focus length of a lens (c.f. Siegman "Lasers" chap. 15)
  •  The ABCDEF matrix from STM1 to ITM is as below
    •  M^=L^pr3itmR^pr3L^pr2pr3L^pr2GrL3F^GrL3L^GrL3GrM3R^GrM3L^GrM3GrM2R^GrM2\hat{M} = \hat{L}_{\mathrm{pr3itm}} \hat{R}_{\mathrm{pr3}} \hat{L}_{\mathrm{pr2pr3}} \hat{L}_{\mathrm{pr2GrL3}} \hat{F}_{\mathrm{GrL3}} \hat{L}_{\mathrm{GrL3GrM3}} \hat{R}_{\mathrm{GrM3}} \hat{L}_{\mathrm{GrM3GrM2}} \hat{R}_{\mathrm{GrM2}}
  • Assuming that the ray vector before shifting is r0=[00]r_0 = \begin{bmatrix} 0 \\ 0 \end{bmatrix}, and that STM1 and STM2 are shifted forward by 1.8um (= 3.6 um /150 V * 75 V) each and no misalignment has occurred, the beam shift on the ITM is equal to M(1,3).
    • M(1,3) ~ 0.1 mm
    • R_e
takahiro.yamamoto - 20:57 Tuesday 09 August 2022 (21743) Print this report

[Ushiba, YamaT]

After this work, 2048Hz and its harmonics (4096Hz and 6144Hz) appears on the control signals of Gr PDH-X and ALS DARM as shown in Fig.1.
Blue and red curves represent before (around 9:00 on 8/8 JST) and after (around 15:00 on 8/8) spectra, respectively.
There is no line on 1024Hz and 3072Hz. So it seems to be a series of 2048Hz.

When we took the red curve, PZT output is 75V (~16384ct) as the offset form DAC output.
And these lines can be removed by changing the DAC output as 0ct (0V).
So the output to POP PZT induces these lines.

Because lines lie just 2048, 4096, and 6144Hz with the frequency resolution as 0.125Hz.
So it seems to be caused from digital issue not from PZT itself, circuits etc.

Tomorrow, Ushiba-kun will check the output signals from Dsub-BNC converter to check that
a cause lie around digital side or PZT side just in case.

We set PZT offset as 0 for today.

Images attached to this comment
takafumi.ushiba - 11:07 Wednesday 10 August 2022 (21747) Print this report

[Yasui, Ushiba]

We checked DAC outputs from corresponding channels and found that there are strange 2048-Hz signals in the DAC output itself (fig1).
During ths measurement, we disconnected the BNC cables between DGS rack and PZT drivers, so it is not likely to be related GND of the drivers.

Images attached to this comment
takahiro.yamamoto - 14:04 Wednesday 10 August 2022 (21749) Print this report

It comes from the digital AI process between user models and IOP models.

On real-time system, default implementation of the digital AI process is 0-padding way.
FIg.1 shows a simulated signal on the 65kHz IOP model when a 2kHz user model outputs 10000ct offset.
We can see a periodical (2kHz) glitch like signals.
By decreasing offset value on the user model, an amplitude of the glitch like signal also decrease (See also Fig.2 with 1000ct offset).
Peak amplitude of these glitches seems to become ~27% larger than the given DC offset value.

We have two choices to avoid this noise.

One is to change the model rate from 2kHz to 16kHz.
A simulated signal on the IOP model when 16kHz user model outputs 10000ct offset is shown in Fig.3.
In this case, 16kHz noise is induced but peak amplitude can be suppressed as ~4% of the given DC offset value.
Thanks to the response of cavity, PZT, etc. noise around 16kHz may be practically negligible.

Another choice is to use "no_zero_pad" option on the user model.
In this case, we can completely remove this kind of noise. But CPU load increase.
Fortunately, CPU load is not so serious because k1ial is now very small model.
(It can be used on very large model e.g k1lsc, k1asc, and so on.)

I haven't used this option and seen it also on LIGO's models yet.
So we should test it if we choose the latter choice.

Images attached to this comment
MIF (ITF Control)
masayuki.nakano - 14:56 Monday 08 August 2022 (21729) Print this report
Comment to ALS DARM was locked (21659)

OMG!! I missed a really great news!!!

Congrats!!!
I was really surprised by your speed to succeed in the type-A control, and I'm pretty sure that you will lock FPMI and be ready for the noise hunting!!

Go KAGRA, go!!

FCL (Air)
shinji.miyoki - 11:13 Monday 08 August 2022 (21728) Print this report
BS-C cooler stopped arpund 9:30

Hayakawa-kun found BS-C cooler stopped around 9:30(It worked around 8:00). The reason could be the reduction of water flow because of the cryo-cooler operation for IYC duct shields on Friday.

So, we should make a set of checks when we operated something that requires water in the corner station because the continuous operation of the Daikin coolers in the corner station is so sensitive to the amount of water flow.

CRY (Cryostat IY)
shinji.miyoki - 10:58 Monday 08 August 2022 (21727) Print this report
Comment to Turned (21710)

I turned off 2 of 10 FFUs in the IYC clean booth. 8 are now operated.

Search Help
×

Warning

×